# VL-ICL Bench: The Devil in the Details of Benchmarking Multimodal In-Context Learning

### v1

@@ 6 Conclusion
  • Section 1: Implementation and Evaluation Details
  • Section 2: Further Experimental Analysis
    Subsection 2.1: Results of Scaling to Many Shots (Figure 10).
    Subsection 2.2: Results of Chain-of-Thought Prompting (Figure 11).
    Subsection 2.3: Results of Repeating the Support Set (Figure 12).
    Subsection 2.4: Influence of Instruction Fine-tuning to ICL (Figure 13).
    Subsection 2.5: Different Levels of Task Descriptions (Figure 14).
  • Section 3: Complete Results
    Raw results of all figures in the main text and appendix from Table 3 to 50.
1 Implementation and Evaluation Details
@@
Prompt to judge image generation for Fast MiniImageNet and CobSAT dataset User: Decide whether the image contains the following concept: {GT}. Answer with ’yes’ or ’no’.
@@
Prompt to judge the answer for Vizwiz VQA. User: Based on the image and question, decide whether the predicted answer has the same meaning as the ground truth. Answer with ’yes’ or ’no’. Question: {Question} Predicted answer: {Prediction} Ground Truth: {GT}
Prompt to rate the quality of COCO captioning (Main text, section 2.2) User: Given the following image, you are to evaluate the provided generated caption based on its relevance, accuracy, completeness, and creativity in describing the image. Rate the caption on a scale from 1 to 10, where 10 represents an exceptional description that accurately and completely reflects the image’s content, and 1 represents a poor description that does not accurately describe the image. Generated Caption: {Prediction} Ground Truth Caption: {GT} Consider the following criteria for your rating: 1 (Very Poor): The caption does not correspond to the image’s content, providing incorrect information or irrelevant descriptions. It misses essential elements and may introduce non-existent aspects. 3 (Poor): The caption only slightly relates to the image, missing significant details or containing inaccuracies. It acknowledges some elements of the image but overlooks key aspects. 5 (Fair): The caption provides a basic description of the image but lacks depth and detail. It captures main elements but misses subtleties and may lack creativity or precision. 7 (Good): The caption accurately describes the main elements of the image, with some attention to detail and creativity. Minor inaccuracies or omissions may be present, but the overall description is sound. 8 (Very Good): The caption provides a detailed and accurate description of the image, with good creativity and insight. It captures both essential and minor elements, offering a well-rounded depiction. 9 (Excellent): The caption delivers an accurate, detailed, and insightful description, demonstrating high creativity and a deep understanding of the image. It covers all relevant details, enhancing the viewer’s perception. 10 (Exceptional): The caption offers a flawless description, providing comprehensive, accurate, and highly creative insights. It perfectly aligns with the image’s content, capturing nuances and offering an enhanced perspective. Please provide your rating. You should ONLY output the score number.
2 Further Analysis
2.1 Scaling to More Shots
@@
To examine the maximum number of shots the models can handle and whether the model can still benefit from more shots, we further increase the support set size to 16, 32, and 64 shots. We choose three models for this experiment: OpenFlamingo 9B [Awadalla et al., 2023], IDEFICS-9B-Instruct [Laurençon et al., 2023], and InternLM-XComposer2 [Zhang et al., 2023]. These models were selected because each image they process translates to fewer tokens (Table 2), ensuring that they do not exceed the maximum context length when evaluated with 64 shots. IDEFICS-9B-Instruct demonstrates a better scaling capability compared to other models in most of the datasets. Besides, while InternLM-XComposer2 has strong performance in a low-shot regime, the performance quickly decreases with many shots. This may be due to the mismatch between training (4096) and testing (2048) context length (Table 2) where the extrapolation of context length has been a well-known challenging task [Press et al., 2022, Liu et al., 2024b].
@@
2.2 Chain-of-Thought Prompting
To investigate whether there is any strategy that can enhance in-context learning, one straightforward method is Chain-of-Thought (CoT) prompting [Wei et al., 2022]. CoT prompts the model to articulate its reasoning process concerning latent variables from the support set, potentially improving its learning and inference capabilities. We experiment with Qwen-VL-Chat [Bai et al., 2023] and InternLM-XComposer2 [Zhang et al., 2023] that have state-of-the-art LLMs with strong reasoning ability. Below is the specific prompt we use. [CoT Prompt]: Let’s first think step by step and analyze the relationship between the given few-shot question-answer pairs. Give reasoning rationales. User: [Task Description][Support Set][Query][CoT Prompt] VLLMs: [Generated rationals] User: [Task Description][Support Set][Query][Generated rationals] VLLMs: Prediction
@@ 2.3 Repeating Support Set
2.4 Influence of Instruction Fine-tuning
@@
2.5 Different Levels of Task Description Details
@@
The task descriptions that we use for the different datasets are as follows:
Fast Open-Ended MiniImageNet Detailed: Induce the concept from the in-context examples. Answer the question with a single word or phase. Concise: Answer the question with a single word or phase.
CLEVR Count Induction Detailed: The image contains objects of different shapes, colors, sizes and materials. The question describes the attribute and its value. You need to find all objects within the image that satisfy the condition. You should induce what operation to use according to the results of the in-context examples and then calculate the result. Concise: Find objects of the given type, induce what operation to use and calculate the result.
Operator Induction Detailed: The image contains two digit numbers and a ? representing the mathematical operator. Induce the mathematical operator (addition, multiplication, minus) according to the results of the in-context examples and calculate the result. Concise: Induce the mathematical operator and calculate the result.
TextOCR Detailed: An image will be provided where a red box is drawn around the text of interest. Answer with the text inside the red box. Ensure that the transcription is precise, reflecting the exact characters, including letters, numbers, symbols. Concise: Answer with the text inside the red box.
@@
3 Complete Results
In this section, we show the raw results of the figures in the main text in Section 3.1 to 3.3, and results of supplementary materials in Section 3.4.
3.1 Initial Analysis of VQA and Image Captioning
@@
3.2 Main Results
We present the main results from Table 12 to 24.

### v2

@@ 6 Discussion
Table of Content
  • Section A: Implementation and Evaluation Details
  • Section B: Further Experimental Analysis
    Subsection B.1: Full Main Results (Figure 8).
    Subsection B.2: Results of Scaling to Many Shots (Figure 9).
    Subsection B.3: Results of Chain-of-Thought Prompting (Figure 10).
    Subsection B.4: Results of Repeating the Support Set (Figure 11).
    Subsection B.5: Influence of Instruction Fine-tuning to ICL (Figure 12).
    Subsection B.6: Different Levels of Task Descriptions (Figure 13).
    Subsection B.7: ICL Efficiency vs Zero-Shot Performance (Figure 14).
    Subsection B.8: Self-Extend Analysis (Figure 15).
    Subsection B.9: Average Performance over Time (Figure 16).
  • Section C: Qualitative Results
  • Section D: Complete Results
    Raw results of all figures in main paper and supplementary materials from Table 7 to 67.
Appendix A Implementation and Evaluation Details
@@
Phi3-Vision MLP AnyRes - 128K
LongVA MLP UniRes - 200K+
IDEFICS2 MLP 64 - 32K
LLaVA-OneVision MLP AnyRes - 128K
InternLM-X2d5 Partial-LoRA AnyRes 24K 96K
@@
[backgroundcolor=gray!20] Prompt to judge image generation for Fast MiniImageNet and CobSAT dataset
User: Decide whether the image contains the following concept: {GT}. Answer with ’yes’ or ’no’.
{mdframed}
[backgroundcolor=gray!20] Prompt to judge the answer for Vizwiz VQA.
@@
User: Based on the image and question, decide whether the predicted answer has the same meaning as the ground truth. Answer with ’yes’ or ’no’. Question: {Question} Predicted answer: {Prediction} Ground Truth: {GT}
{mdframed}
[backgroundcolor=gray!20] Prompt to rate the quality of COCO captioning (Main text, section 2.2)
User: Given the following image, you are to evaluate the provided generated caption based on its relevance, accuracy, completeness, and creativity in describing the image. Rate the caption on a scale from 1 to 10, where 10 represents an exceptional description that accurately and completely reflects the image’s content, and 1 represents a poor description that does not accurately describe the image.
Generated Caption: {Prediction}
Ground Truth Caption: {GT}
Consider the following criteria for your rating:
1 (Very Poor): The caption does not correspond to the image’s content, providing incorrect information or irrelevant descriptions. It misses essential elements and may introduce non-existent aspects.
3 (Poor): The caption only slightly relates to the image, missing significant details or containing inaccuracies. It acknowledges some elements of the image but overlooks key aspects.
5 (Fair): The caption provides a basic description of the image but lacks depth and detail. It captures main elements but misses subtleties and may lack creativity or precision.
7 (Good): The caption accurately describes the main elements of the image, with some attention to detail and creativity. Minor inaccuracies or omissions may be present, but the overall description is sound.
8 (Very Good): The caption provides a detailed and accurate description of the image, with good creativity and insight. It captures both essential and minor elements, offering a well-rounded depiction.
9 (Excellent): The caption delivers an accurate, detailed, and insightful description, demonstrating high creativity and a deep understanding of the image. It covers all relevant details, enhancing the viewer’s perception.
10 (Exceptional): The caption offers a flawless description, providing comprehensive, accurate, and highly creative insights. It perfectly aligns with the image’s content, capturing nuances and offering an enhanced perspective.
Please provide your rating. You should ONLY output the score number.
Appendix B Further Analysis
B.1 Full VL-ICL Bench  Results
We show the main results across all datasets in Fig. 8.
Figure 8: VL-ICL Bench results. Top: Image-to-Text. Bottom: Text-to-Image tasks.
B.2 Scaling to More Shots
@@
To examine the maximum number of shots the models can handle and whether the model can still benefit from more shots, we further increase the support set size to 16, 32, and 64 shots. We choose three models for this experiment: OpenFlamingo 9B [Awadalla et al., 2023], IDEFICS-9B-Instruct [Laurençon et al., 2023], InternLM-XComposer2 [Zhang et al., 2023], LongVA [Zhang et al., 2024b], and Mantis-Idefics2 [Jiang et al., 2024a]. These models were selected because each image they process translates to fewer tokens (Table 6) or they have a longer context window, ensuring that they do not exceed the maximum context length when evaluated with 64 shots. IDEFICS-9B-Instruct demonstrates a better scaling capability compared to other models in most of the datasets. Besides, while InternLM-XComposer2 has strong performance in a low-shot regime, the performance quickly decreases with many shots. This may be due to the mismatch between training (4096) and testing (2048) context length (Table 6) where the extrapolation of context length has been a well-known challenging task [Press et al., 2022, Liu et al., 2024b].
@@
B.3 Chain-of-Thought Prompting
To investigate whether there is any strategy that can enhance in-context learning, one straightforward method is Chain-of-Thought (CoT) prompting [Wei et al., 2022b]. CoT prompts the model to articulate its reasoning process concerning latent variables from the support set, potentially improving its learning and inference capabilities. We experiment with Qwen-VL-Chat [Bai et al., 2023] and InternLM-XComposer2 [Zhang et al., 2023] that have state-of-the-art LLMs with strong reasoning ability. Below is the specific prompt we use. {mdframed}[backgroundcolor=gray!20] [CoT Prompt]: Let’s first think step by step and analyze the relationship between the given few-shot question-answer pairs. Give reasoning rationales.
User: [Task Description][Support Set][Query][CoT Prompt]
VLLMs: [Generated rationals]
User: [Task Description][Support Set][Query][Generated rationals]
VLLMs: Prediction
@@ B.4 Repeating Support Set
B.5 Influence of Instruction Fine-tuning
@@
B.6 Different Levels of Task Description Details
@@
The task descriptions that we use for the different datasets are as follows: {mdframed}[backgroundcolor=gray!20] Fast Open-Ended MiniImageNet
Detailed: Induce the concept from the in-context examples. Answer the question with a single word or phase.
Concise: Answer the question with a single word or phase.
{mdframed}
[backgroundcolor=gray!20] CLEVR Count Induction
Detailed: The image contains objects of different shapes, colors, sizes and materials. The question describes the attribute and its value. You need to find all objects within the image that satisfy the condition. You should induce what operation to use according to the results of the in-context examples and then calculate the result.
Concise: Find objects of the given type, induce what operation to use and calculate the result.
{mdframed}
[backgroundcolor=gray!20] Operator Induction
Detailed: The image contains two digit numbers and a ? representing the mathematical operator. Induce the mathematical operator (addition, multiplication, minus) according to the results of the in-context examples and calculate the result.
Concise: Induce the mathematical operator and calculate the result.
{mdframed}
[backgroundcolor=gray!20] TextOCR
Detailed: An image will be provided where a red box is drawn around the text of interest. Answer with the text inside the red box. Ensure that the transcription is precise, reflecting the exact characters, including letters, numbers, symbols.
Concise: Answer with the text inside the red box.
@@
B.7 ICL Efficiency vs Zero-Shot Performance
We analyse the in-context learning efficiency and compare it against the zero-shot performance. We define ICL efficiency as the ability to improve the performance after seeing a few examples. We calculate its value as a proportion of the area between the zero-shot curve and the few-shot curve, over the total available area (0-100% accuracy, max. number of shots). The results are shown in Fig. 14, where we only show models with non-negative efficiency. We also show the pareto front line that describes the constraint on the best possible models: zero-shot performance + ICL efficiency = 100%. The results indicate that there is a good variability between different models in how efficiently they can learn from the provided support examples.
Figure 14: ICL efficiency vs zero-shot performance results. Top: Image-to-Text. Bottom: Text-to-Image tasks.
B.8 Self-Extend Analysis
We analyse the impact of using self-extend mechanism to improve the ability to handle longer context. The results are shown in Figure 15 and were described earlier in the main text.
Figure 15: Comparison of models without self-extend (solid line, circle markers) and with self-extend (dashed line, x markers).
B.9 Average Performance over Time
We analyse in Figure 16 how the average accuracy and ICL efficiency of image-to-text and text-to-image models evolved over time on our tasks. We can see that the performance as well as ICL efficiency have been continuously improving even though there are some outliers (e.g. GPT4V) that were released earlier and obtained some of the stronger performances. This shows that our benchmark can effectively assess the ICL capability of the VLMs.
Figure 16: Average accuracy and ICL eff. in terms of when the model was released.
Appendix C Qualitative Analysis
We also include a qualitative analysis on selected cases, where we analyse the impact of using more support examples on the quality of the output. We analyse text-to-image tasks in Fig. 17, using Emu2-Gen model. For text-to-image MiniImageNet the model should learn from the support examples that the artificial names slation and shously correspond to a lion and a school bus, respectively. Emu2-Gen is able to do it to a certain extent, but may get confused by additional support examples as more support examples are not necessarily helpful. In CoBSAT, the support set induces that the animal should have glacier and desert background. With no support examples the model only displays the animal, but with more support examples it learns that it should use glacier background. In the second example, the model is able to capture that it should use desert background, but is less successful in showing the required animal – zebra. The quality of the generated images is not necessarily better with more support examples.
CLEVR  In many cases the model rephrases the question, while in others it talks that e.g. the described object is present. Such behaviour is more common with fewer or no support examples. With more support examples the model learns to predict a count but gets incorrect answer. It can be because some objects are more difficult to recognize, e.g. if one partially covers another.
Appendix D Complete Results
In this section, we show the raw results of the figures in the main text in Section D.1 to 31, and results of supplementary materials in Section D.4.
D.1 Initial Analysis of VQA and Image Captioning
@@
D.2 Main Results
@@
VILA-2.7B 0.00 $\pm$ 0.00 5.67 $\pm$ 0.24 50.33 $\pm$ 1.18 45.00 $\pm$ 1.08 42.00 $\pm$ 1.78
VILA-7B 0.00 $\pm$ 0.00 28.83 $\pm$ 1.43 38.17 $\pm$ 2.72 37.67 $\pm$ 3.68 37.33 $\pm$ 1.18
Mantis-Idefics2 0.00 $\pm$ 0.00 2.50 $\pm$ 1.08 72.83 $\pm$ 2.39 84.33 $\pm$ 2.09 76.00 $\pm$ 0.41
Phi3-Vision 0.00 $\pm$ 0.00 12.67 $\pm$ 0.85 50.00 $\pm$ 1.08 39.00 $\pm$ 2.48 40.17 $\pm$ 1.25
LongVA-7B 0.00 $\pm$ 0.00 22.33 $\pm$ 0.24 52.67 $\pm$ 3.09 48.83 $\pm$ 1.43 49.50 $\pm$ 1.08
LLaVA-OneVision-72B 0.00 $\pm$ 0.00 12.67 $\pm$ 0.94 88.00 $\pm$ 0.82 98.67 $\pm$ 0.47 98.33 $\pm$ 0.62
InternLM-X2d5 0.00 $\pm$ 0.00 12.00 $\pm$ 0.00 82.50 $\pm$ 1.22 88.17 $\pm$ 1.70 91.00 $\pm$ 1.47
@@
Mantis-Idefics2 18.00 $\pm$ 0.00 27.00 $\pm$ 1.08 49.33 $\pm$ 1.25 16.67 $\pm$ 1.70 17.83 $\pm$ 2.05
idefics2-8b 25.50 $\pm$ 0.00 24.17 $\pm$ 1.25 57.83 $\pm$ 1.31 79.00 $\pm$ 2.55 79.83 $\pm$ 1.25
InternLM-X2d5 18.50 $\pm$ 0.00 41.67 $\pm$ 1.31 69.83 $\pm$ 0.24 83.50 $\pm$ 0.82 84.50 $\pm$ 0.71
LLaVA-OneVision-7B 24.50 $\pm$ 0.00 54.17 $\pm$ 0.85 68.17 $\pm$ 1.31 67.00 $\pm$ 1.08 67.17 $\pm$ 3.42
@@
VILA-2.7B 16.67 $\pm$ 0.00 12.78 $\pm$ 2.08 $1$$\pm$ 3.14 11.67 $\pm$ 1.36 11.67 $\pm$ 2.72
VILA-7B 28.33 $\pm$ 0.00 $1$$\pm$ 4.37 6.67 $\pm$ 3.60 7.78 $\pm$ 0.78 8.33 $\pm$ 2.72
Phi3-Vision 0.00 $\pm$ 0.00 42.22 $\pm$ 1.57 45.56 $\pm$ 0.79 51.67 $\pm$ 1.36 54.44 $\pm$ 2.83
LongVA-7B 18.33 $\pm$ 0.00 8.33 $\pm$ 3.60 6.11 $\pm$ 2.83 5.56 $\pm$ 2.83 3.89 $\pm$ 2.83
Mantis-Idefics2 16.67 $\pm$ 0.00 12.78 $\pm$ 2.08 8.89 $\pm$ 2.08 14.44 $\pm$ 4.16 16.11 $\pm$ 0.79
LLaVA-OneVision-72B $3$$\pm$ 0.00 75.56 $\pm$ 3.42 60.56 $\pm$ 6.71 68.89 $\pm$ 1.57 57.78 $\pm$ 2.08
InternLM-X2d5 36.67 $\pm$ 0.00 40.00 $\pm$ 3.60 41.67 $\pm$ 2.72 35.00 $\pm$ 2.36 35.00 $\pm$ 4.08
@@
VILA-2.7B 0.50 $\pm$ 0.00 6.33 $\pm$ 1.18 14.83 $\pm$ 2.39 16.67 $\pm$ 0.62 18.50 $\pm$ 1.22
VILA-7B 28.00 $\pm$ 0.00 6.17 $\pm$ 1.03 22.17 $\pm$ 0.24 26.83 $\pm$ 0.47 30.17 $\pm$ 1.03
Phi3-Vision 8.00 $\pm$ 0.00 32.17 $\pm$ 2.25 38.17 $\pm$ 0.85 39.67 $\pm$ 0.24 41.50 $\pm$ 1.78
LongVA-7B 26.00 $\pm$ 0.00 3.50 $\pm$ 0.82 11.17 $\pm$ 0.85 17.17 $\pm$ 1.65 17.50 $\pm$ 2.12
Mantis-Idefics2 21.00 $\pm$ 0.00 13.83 $\pm$ 1.18 25.17 $\pm$ 0.62 27.83 $\pm$ 1.25 27.67 $\pm$ 0.47
LLaVA-OneVision-72B 48.50 $\pm$ 0.00 50.67 $\pm$ 1.70 51.17 $\pm$ 0.24 50.67 $\pm$ 2.09 51.67 $\pm$ 0.62
InternLM-X2d5 32.00 $\pm$ 0.00 36.50 $\pm$ 1.41 37.00 $\pm$ 0.41 42.50 $\pm$ 0.71 27.02 $\pm$ 0.24
@@
VILA-2.7B 0.00 $\pm$ 0.00 30.67 $\pm$ 2.09 29.17 $\pm$ 0.85 29.17 $\pm$ 1.03 29.83 $\pm$ 0.24
VILA-7B 3.50 $\pm$ 0.00 34.00 $\pm$ 2.86 32.67 $\pm$ 1.43 31.50 $\pm$ 2.16 34.33 $\pm$ 2.39
Phi3-Vision 0.00 $\pm$ 0.00 34.50 $\pm$ 2.55 25.50 $\pm$ 2.45 17.67 $\pm$ 2.09 18.67 $\pm$ 0.85
LongVA-7B 0.00 $\pm$ 0.00 29.17 $\pm$ 1.31 22.83 $\pm$ 2.39 20.00 $\pm$ 1.08 23.83 $\pm$ 3.47
Mantis-Idefics2 31.45 $\pm$ 2.18 40.17 $\pm$ 3.12 32.00 $\pm$ 0.71 24.50 $\pm$ 1.08 21.67 $\pm$ 2.05
LLaVA-OneVision-72B 0.50 $\pm$ 0.00 37.17 $\pm$ 2.01 32.00 $\pm$ 0.71 31.83 $\pm$ 1.84 42.33 $\pm$ 0.85
InternLM-X2d5 6.00 $\pm$ 0.00 63.17 $\pm$ 1.25 58.50 $\pm$ 2.86 56.00 $\pm$ 2.68 53.17 $\pm$ 2.49
@@
Table 21: Results of different models on Fast Matching MiniImageNet dataset (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 5-Shot
OpenFlamingo-9B 0.00 $\pm$ 0.00 18.25 $\pm$ 1.32 19.28 $\pm$ 5.17 26.33 $\pm$ 3.02 29.11 $\pm$ 1.30
IDEFICS-9B 0.00 $\pm$ 0.00 0.08 $\pm$ 0.12 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
IDEFICS-80B 0.00 $\pm$ 0.00 21.58 $\pm$ 3.70 28.30 $\pm$ 1.94 23.16 $\pm$ 1.45 25.58 $\pm$ 2.13
Otter 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
IDEFICS2-8B 0.00 $\pm$ 0.00 0.58 $\pm$ 0.12 8.33 $\pm$ 1.16 17.33 $\pm$ 0.92 25.25 $\pm$ 1.74
Mantis-Idefics2 0.00 $\pm$ 0.00 51.08 $\pm$ 3.44 49.25 $\pm$ 1.59 47.08 $\pm$ 0.72 47.17 $\pm$ 0.47
InternLM-X2 0.00 $\pm$ 0.00 49.92 $\pm$ 0.12 49.92 $\pm$ 0.12 0.92 $\pm$ 0.31 0.83 $\pm$ 0.24
Qwen-VL-Chat 0.00 $\pm$ 0.00 0.33 $\pm$ 0.24 0.00 $\pm$ 0.00 0.08 $\pm$ 0.12 0.08 $\pm$ 0.12
Phi3-Vision 0.00 $\pm$ 0.00 45.92 $\pm$ 0.85 46.08 $\pm$ 1.48 47.58 $\pm$ 1.23 50.75 $\pm$ 1.43
LLaVA-Next-7B 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
LongVA-7B 0.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00
VILA-7B 0.00 $\pm$ 0.00 49.67 $\pm$ 0.31 49.75 $\pm$ 0.35 49.08 $\pm$ 0.77 49.92 $\pm$ 0.24
Emu2-Chat 0.00 $\pm$ 0.00 32.59 $\pm$ 0.58 36.80 $\pm$ 1.80 37.62 $\pm$ 0.93 40.25 $\pm$ 2.43
InternLM-X2d5 0.00 $\pm$ 0.00 9.25 $\pm$ 0.00 46.75 $\pm$ 0.00 31.50 $\pm$ 0.00 8.00 $\pm$ 0.00
LLaVA-OneVision-72B 0.00 $\pm$ 0.00 28.50 $\pm$ 2.07 58.08 $\pm$ 0.92 68.83 $\pm$ 0.31 70.33 $\pm$ 0.66
GPT4V 0.00 2.00 48.75 58.25 58.00
@@
VILA-2.7B 13.33 $\pm$ 0.00 10.56 $\pm$ 0.79 10.56 $\pm$ 0.79 $1$$\pm$ 2.08 11.67 $\pm$ 1.36
VILA-7B 28.33 $\pm$ 0.00 6.11 $\pm$ 2.08 11.67 $\pm$ 3.60 12.22 $\pm$ 2.83 13.33 $\pm$ 2.36
Phi3-Vision 0.00 $\pm$ 0.00 16.67 $\pm$ 2.72 21.67 $\pm$ 1.36 15.56 $\pm$ 1.57 27.78 $\pm$ 3.14
LongVA-7B 10.00 $\pm$ 0.00 5.00 $\pm$ 1.36 4.44 $\pm$ 0.79 4.45 $\pm$ 3.14 3.89 $\pm$ 4.37
Mantis-Idefics2 30.00 $\pm$ 0.00 5.56 $\pm$ 3.42 4.44 $\pm$ 0.79 6.67 $\pm$ 1.36 7.22 $\pm$ 0.78
LLaVA-OneVision-72B 38.33 $\pm$ 0.00 42.78 $\pm$ 2.08 41.11 $\pm$ 2.08 36.67 $\pm$ 2.72 47.78 $\pm$ 10.03
InternLM-X2d5 38.33 $\pm$ 0.00 32.22 $\pm$ 3.43 31.67 $\pm$ 3.60 33.89 $\pm$ 2.08 21.11 $\pm$ 2.08
@@
VILA-2.7B 59.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00
VILA-7B 77.25 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00
Mantis-Idefics2 66.00 $\pm$ 0.00 54.17 $\pm$ 0.66 55.25 $\pm$ 1.47 50.25 $\pm$ 0.20 49.75 $\pm$ 0.35
Phi3-Vision 51.75 $\pm$ 0.00 50.08 $\pm$ 0.12 50.75 $\pm$ 0.20 52.17 $\pm$ 1.25 51.25 $\pm$ 0.71
LongVA-7B 58.25 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00
LLaVA-OneVision-72B 93.00 $\pm$ 0.00 87.12 $\pm$ 0.12 93.12 $\pm$ 1.38 91.62 $\pm$ 0.62 92.62 $\pm$ 0.12
InternLM-X2d5 83.75 $\pm$ 0.00 74.92 $\pm$ 0.62 84.67 $\pm$ 0.92 69.17 $\pm$ 1.01 50.17 $\pm$ 0.12
@@
Anole-7B 0.50 $\pm$ 0.41 11.00 $\pm$ 2.86 7.00 $\pm$ 0.71 0.17 $\pm$ 0.24 0.17 $\pm$ 0.24
Table 28: Results of different models on Fast Counting (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 5-Shot
GILL 34.17 $\pm$ 2.36 8.33 $\pm$ 4.25 3.33 $\pm$ 2.36 3.33 $\pm$ 1.18 7.50 $\pm$ 2.04
SEED-LLaMA-8B 34.17 $\pm$ 2.36 50.00 $\pm$ 6.12 47.50 $\pm$ 4.08 45.00 $\pm$ 2.04 56.67 $\pm$ 6.24
SEED-LLaMA-14B 41.26 $\pm$ 1.25 42.72 $\pm$ 5.14 37.50 $\pm$ 2.36 51.57 $\pm$ 3.25 37.55 $\pm$ 5.54
Emu1-Gen 33.19 $\pm$ 4.27 39.55 $\pm$ 4.51 37.56 $\pm$ 3.17 47.81 $\pm$ 3.94 45.52 $\pm$ 3.57
Emu2-Gen 44.17 $\pm$ 4.32 50.80 $\pm$ 2.12 47.13 $\pm$ 3.98 49.25 $\pm$ 4.04 59.25 $\pm$ 5.23
Table 29: Results of different models on Fast Attribute Matching (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 5-Shot
GILL 20.50 $\pm$ 1.08 6.50 $\pm$ 0.41 4.67 $\pm$ 1.43 6.17 $\pm$ 1.31 7.00 $\pm$ 1.47
SEED-LLaMA-8B 21.00 $\pm$ 0.41 31.17 $\pm$ 1.43 34.33 $\pm$ 0.62 32.17 $\pm$ 1.03 32.17 $\pm$ 1.03
SEED-LLaMA-14B 22.36 $\pm$ 0.28 26.35 $\pm$ 1.93 33.57 $\pm$ 2.01 32.00 $\pm$ 1.94 35.59 $\pm$ 3.09
Emu1-Gen 23.47 $\pm$ 3.98 24.11 $\pm$ 6.26 27.58 $\pm$ 2.47 29.14 $\pm$ 5.98 24.27 $\pm$ 1.86
emu2-gen 26.50 $\pm$ 1.78 18.67 $\pm$ 1.03 23.17 $\pm$ 1.43 34.00 $\pm$ 2.27 28.00 $\pm$ 1.08
D.3 Additional Analysis